OpenStudy (anonymous):

substitution, 3x-2y=0 2x-y=-1

OpenStudy (anonymous):

3x - 2y = 0 2x - y = -1 to make it easier, multiply 2 in the 2nd equation. 3x - 2y = 0 4x - 2y = -2 Subtract both so you can find x - x = 2 x = -2 substitute x to one of the given equations to find y. 1st eq 3(-2) - 2y = 0 -6 - 2y = 0 -2y = 6 y = -3 2nd eq 2(-2) - y = -1 -y = 3 y = -3 therefore, x = -2 and y = -3 hope this helps! ;)